OpenStudy (anonymous):

Carol spends 17 hours in a 2-week period practicing her culinary skills. How many hours does she practice in 5 weeks?

OpenStudy (lifeisadangerousgame):

Well, every two weeks its 17 hours correct?

OpenStudy (anonymous):

Yeah

OpenStudy (lifeisadangerousgame):

We have 5 weeks to work with. Now every two weeks we can add 17. So the first two +17 and the second two +17. That's 34 hours. We have one week left to work with. Now to get this we take 17 and divide it, since its only one week and not two. That's 8.5. SO if I'm doing this right, its 42.5 hours.

OpenStudy (ranga):

This is a direct proportions problem: 17/2 = x/5. Solve for x.

OpenStudy (anonymous):

is it 85?

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

x/ 5 = 17/2 Isolate x. So multiply both sides by 5 x = 17/2 x 5 = ?

OpenStudy (anonymous):

425?

OpenStudy (ranga):

Yes! Same as the other person's answer before.

OpenStudy (anonymous):

Oh thanks.

OpenStudy (anonymous):

don't forget the decimal.

OpenStudy (ranga):

yeah, it is 42.5

OpenStudy (lifeisadangerousgame):

Yea don't forget the decimal cupcake_queen, it's important! ^_^

OpenStudy (anonymous):

Loll oh okay . thanks yaall

OpenStudy (ranga):

You are welcome. Was that just a typo? (425 instead of 42.5?)

OpenStudy (anonymous):

yeah , I just noticed that xD